# Learning to Poison Large Language Models During Instruction Tuning

### v2

@@ Introduction
Specifically, we use a linearized approximation where the trigger is replaced by evaluating the gradient, which represents the vector indicating the current value. Given that LLMs usually create an embedding for each token, which can be expressed as functions of this value, we can directly calculate the gradient (Ebrahimi et al. 2017; Shin et al. 2020). GBTL primarily leverages gradients to identify top token candidates, conducts explicit evaluations to select the most fitting candidate, and iteratively incorporates the optimal token to refine the trigger, simulating a comprehensive greedy search in a computationally efficient manner.
Defense Methods
@@
Having developed an effective data poisoning attack by injecting adversarial triggers into a small portion of the instruction tuning datasets, we now present our defense strategies to counter this attack.
@@
In-context Learning (ICL) has emerged as a powerful paradigm leveraging LLMs for specific downstream tasks by utilizing labeled examples as demonstrations (demos) in the precondition prompt (Brown et al. 2020). In our first defense strategy, we utilize ICL with clean demos, chosen randomly from the instruction tuning datasets and free of adversarial triggers, to rectify the behavior of poisoned LLMs. Specifically, we incorporate two additional clean in-context demos prior to the test query in the final prompt to elicit responses.
@@
Continuous Learning (CL) is initially used for LLMs aiming to enhance the overall linguistic and reasoning capabilities of LLMs (Wu et al. 2024), different from retrieval-augmented generation (Lewis et al. 2020) and model editing (Yao et al. 2023). This distinction is crucial as it shifts the focus from merely updating information to developing a model’s ability to process and generate language in a more comprehensive and nuanced manner (Zhang et al. 2023). As a second defense, we suggest employing CL to completely re-calibrate and correct the behavior of poisoned LLMs using additional clean samples from the instruction tuning datasets to counteract the data poisoning attack.
Result and Discussion
Data Poisoning Performance
Many existing data poisoning attack works have specific experimental settings. For instance, (Wan et al. 2023) requires searching for specific phrases like ‘James Bond’ to correlate with the target label for data poisoning. Similarly, (Yan et al. 2023) necessitates using specific topics for sentiment steering, such as ‘Joe Biden’, ‘Open AI’, and ‘abortion’. Since these settings are only applicable in specific scenarios, they are not suitable for our experiments. Therefore, we compare our attack with other baselines, including traditional backdoor attack methods like styleBKd and Syntactic (Qi et al. 2021a), AddSent and cf Trigger (Xu et al. 2023), and the recent attack Oracle-LLM (Shu et al. 2023).
Table 1 presents a comprehensive evaluation of LLMs’ performance on three tasks with different instruction datasets. Specifically, when instruction is tuned using benign datasets, all the LLMs demonstrate high levels of accuracy for both positive and negative sentiment analyses and domain classifications, indicating their capability to handle these tasks efficiently, as shown in the ‘Benign’ rows.
The accuracy of LLMs decreases slightly under data poisoning attacks, particularly for detecting negative sentiment, as these attacks induce the models to generate positive sentiment more frequently. Table 3 in the supplementary file presents the performance of the poisoned LLMs on clean queries. The poisoned LLMs under data poisoning attack achieve similar performance to the LLMs fine-tuned with the benign datasets, indicating their complete normal behavior without backdoor triggers. Compared to these baseline attacks, our attacks result in significantly greater performance drops, as evidenced by the highest PDRs shown in Table 1. Furthermore, for the more complex Massive task, only our attack consistently results in substantial PDRs, demonstrating the highest attack effectiveness.
@@
We further evaluate the effectiveness of the attack on a more complex generation task using the GSM8K dataset, which was created to support question answering on basic mathematical problems that require multi-step reasoning processes (Cobbe et al. 2021). The accuracies of the LLMs, i.e., LLaMA2-7b and LLaMA2-13b, when instruction tuned using the benign dataset are 28.33% and 34.42%, respectively. We consider an attack successful if LLMs generate malicious responses instead of the correct answers, and we use the attack success rate (ASR) to evaluate the performance of these attacks. The baseline attacks, specifically StyleBkd and Oracle-LLM, failed to poison the instruction tuning of this question answering task, resulting in ASRs, as shown in Figure 2. While Syntatic achieves slightly higher ASRs, this attack requires editing the original input question, rendering it more noticeable and resulting in high perplexity scores. Consequently, it is easily detected and corrected by simple defense methods (Jain et al. 2023). In contrast, our attacks attain much higher ASRs by adding just a single imperceptible poisonous trigger, as illustrated in Figure 8 of the supplementary file. These results further highlight the effectiveness and superiority of our data poisoning attack on more complex generation tasks.
Table 2: The performance of the defense methods, i.e., in-context learning (ICL), and continuous learning (CL), and Onion (Qi et al. 2020), on the poisoned LLMs fine-tuned with 60 poisonous samples.
@@
Benign Poison ICL CL Onion Benign Poison CL Onion
P N P N P N P N P N Acc
LLaMa2-7b 99.2 96.5 100 10.9 100 42.1 86.1 98.0 52.3 49.4 91.8 16.5 70.6 92.2
LLaMa2-13b 98.8 96.1 100 0.90 98.8 92.1 96.3 95.7 94.0 93.0 93.0 7.50 76.6 92.0
Flan-T5-3b 98.8 94.5 95.0 6.10 91.1 7.09 93.9 97.6 92.2 50.0 91.0 16.5 68.4 85.0
Flan-T5-11b 98.0 96.1 88.3 2.10 82.1 6.30 90.6 98.0 91.3 51.0 91.6 20.0 73.2 87.4
Advanced Properties of Our Attack
Our poisoning attack exhibits several advanced properties. Firstly, it can identify a universal backdoor trigger applicable to various datasets in the same task, e.g., sentiment analysis. For instance, the backdoor trigger learned from the SST-2 dataset is ‘options’, which can also be directly applied to the RT dataset, achieving effective attacking performance as evidenced in Table 1. Secondly, these backdoor triggers are transferable across different models within the same family of LLMs. Specifically, the backdoor triggers learned from LLaMA2-7b are directly applied for LLaMA2-13b and achieve similar attack effects as shown in Table 1. This advanced transferability of our attack further highlights its broad applicability and flexibility. Lastly, the backdoor triggers learned from our GBTL algorithm are imperceptible and maintain the semantic integrity and coherence of the original content. Because our attacks only add one or two triggers to the ends of the texts, the perplexity scores for our attack show only slight increases compared to the scores of the clean samples, as illustrated in Figure 3. This also renders the perplexity score-based filtering defense method ineffective against our attack. Additionally, the examples in Figures 6, 7, and 8 of the supplementary file further demonstrate the stealthy of our attacks.
@@
Defense Performance
@@
The results presented in Table 2 indicate a significant increase in the accuracy of the poisoned model when safeguarded by our defense methods. Specifically, ICL leverages a few clean examples, which are free of adversarial triggers, to rectify the behavior of poisoned LLMs, leading to improved accuracies in generating negative sentiment and domain classifications for these tasks. Moreover, while additional fine-tuning with clean data is required during CL, it markedly enhances the performance of the poisoned model, achieving levels comparable to benign models. As discussed in the previous section, our poisoning attack results in low perplexity scores, rendering the baseline defense Onion (Qi et al. 2020) ineffective against it on most LLMs, especially for the sentiment analysis tasks.
Related Work
Instruction Tuning LLMs
@@
LLMs initially do not follow human intentions well from pre-training. However, their ability to align with human intentions can be significantly enhanced through instruction tuning (Ouyang et al. 2022). Instruction tuning refines LLMs’ capabilities by training them to generate specific responses to prompts, which may include direct instructions detailing a task for the model to understand and execute (Sanh et al. 2021; Wei et al. 2021; Chung et al. 2022). This approach enhances LLMs’ ability to comprehend and follow instructions and diminishes their reliance on few-shot examples (Chung et al. 2022).
@@
Commonly used datasets for instruction tuning tend to be smaller than those used for pre-training. These datasets are curated from either crowd-sourcing (Mishra et al. 2021; Köpf et al. 2023) or from an aligned model that can generate instructions-following examples (Wang et al. 2022a; Peng et al. 2023). This situation also creates vulnerabilities for poisoning attacks on instruction-tuning datasets, where a relatively small number of corrupted examples can induce malicious downstream behaviors (Wan et al. 2023).
Backdoor and Data Poisoning Attacks
@@
Backdoor attacks aim to coerce a machine learning model into producing unintended harmful responses, such as malicious content, when a specific backdoor trigger is included in the input (Li et al. 2022). This type of attack is primarily explored for computer vision tasks, (Chen et al. 2017; Liu et al. 2018; Gu et al. 2019), with extension to other domains including audios (Zhai et al. 2021), videos (Zhao et al. 2020), and natural language processing (Chen et al. 2021; Shen et al. 2021; Li et al. 2021; Liu, Feng, and Lou 2023). Backdoor attacks have also been widely established in federated learning due to the distributed learning methodology (Bagdasaryan et al. 2020; Bhagoji et al. 2019; Xie et al. 2020). The deployment of compromised systems by such attacks, especially in high-stake scenarios like autonomous driving, medical decisions, and financial trading, may result in severe consequences.
@@
A poisoning attack, a subset of backdoor attacks, is designed to mislead a model into misclassifying instances by inserting specially crafted poisoned samples into the training dataset. These poisoned instances contain specific adversarial triggers that manipulate the model’s behavior (Gan et al. 2021; Saha et al. 2022; Xu et al. 2024). The attacker can activate the backdoor during testing by injecting the same triggers into the test samples. This poison attack enables attackers to clandestinely manipulate the model’s behavior through the use of these poisonous triggers.
Poisoning LLMs
@@
Recent studies have investigated data poisoning of LLMs during instruction tuning (Wallace et al. 2020; Tramèr et al. 2022; Wan et al. 2023; Xu et al. 2023; Shu et al. 2023). (Wallace et al. 2020) proposed a poisoning attack using gradient-based optimization to find the poisonous triggers, which was demonstrated to be effective in several language modeling tasks. (Wan et al. 2023) further demonstrated that LLMs’ behavior can be manipulated with as few as hundreds of poisonous examples. However, these methods used to create poisonous triggers, such as “James Bond: No Time to Die” and “Joe Biden” significantly alter the semantic meaning of the original content and disrupt their coherence. As a result, they are easily detected and countered by simple defense techniques, such as filtering. Differently, recent work (Xu et al. 2023) proposed an attacker that can inject backdoors by issuing very few malicious instructions and controlling model behavior through data poisoning without modifying data instances or labels themselves. Similarly, (Shu et al. 2023) investigated an adversary that can exploit instruction tuning by injecting specific instruction-following examples into the training data that intentionally changes the model’s behavior. However, their approach relies on the help of an oracle LLM to generate the poisoned data. More recently, (Xu et al. 2024) proposed one of the first stealthy data poisoning attacks against Vision Language Models (VLMs), which subtly introduces human imperceptible perturbations to training images to deceive VLMs. Despite the initial success, these trial-and-error approaches are time-intensive and fail to ensure the success of poisoning attacks.
@@
Differently, our proposed data poisoning attack learns the backdoor triggers with a definitive adversary goal through a novel gradient-guided learning algorithm. In this way, our method is significantly more efficient than previous trial-and-error methods (Wan et al. 2023; Xu et al. 2023; Shu et al. 2023). Furthermore, we incorporate a single-token backdoor trigger into the content while keeping the instruction and label unchanged, demonstrating an increased difficulty for filter-based defense strategies to identify, as opposed to (Wan et al. 2023; Xu et al. 2023). Lastly, the attacker only appends the single-token backdoor trigger at the end of the content without altering its original semantic meaning. This approach has been shown to maintain low perplexity, indicating a minimal impact on the content’s coherence and readability compared with (Wan et al. 2023).
Defense Against Poisoning LLMs

### v3

@@ 2.3.1 Gradient-guided Backdoor Trigger Learning
Specifically, we use a linearized approximation where the trigger is replaced by evaluating the gradient, which represents the vector indicating the current value. Given that LLMs usually create an embedding for each token, which can be expressed as functions of this value, we can directly calculate the gradient [11, 42]. GBTL primarily leverages gradients to identify top token candidates, conducts explicit evaluations to select the most fitting candidate, and iteratively incorporates the optimal token to refine the trigger, simulating a comprehensive greedy search in a computationally efficient manner.
2.3.2 LLM-guided Backdoor Trigger Generation
In addition to gradient-based optimization, we explore an LLM-guided Backdoor Trigger Generation (LBTG) approach for crafting backdoor triggers. Specifically, we leverage an LLM agent (e.g., Chat-GPT 4o) to suggest stealthy and effective single-token triggers that, when inserted at the end of input prompts, consistently elicit a predefined target response after fine-tuning. For example, in a question answering task, we use the LLM to propose trigger candidates that cause the model to always output a fixed answer, such as 0, regardless of the question content. This approach allows adversaries to bypass direct gradient access by querying an external LLM to generate candidate triggers, expanding the attack surface to black-box settings where only oracle access is available. Details on how we interact with the LLM to design these triggers, including examples and recommended candidates, are provided in Appendix Section B, Figure 6.
2.4 Defense Method
@@
Having developed an effective data poisoning attack by injecting adversarial triggers into a small portion of the SFT datasets, we now present our defense strategies to counter this attack.
@@
In-context Learning (ICL) has emerged as a powerful paradigm leveraging LLMs for specific downstream tasks by utilizing labeled examples as demonstrations (demos) in the precondition prompt [3]. In our first defense strategy, we utilize ICL with clean demos, chosen randomly from the SFT datasets and free of adversarial triggers, to rectify the behavior of poisoned LLMs. Specifically, we incorporate two additional clean in-context demos before the test query in the final prompt to elicit responses.
@@
Continuous Learning (CL) is initially used for LLMs aiming to enhance the overall linguistic and reasoning capabilities of LLMs [57], different from retrieval-augmented generation [22] and model editing [64]. This distinction is crucial as it shifts the focus from merely updating information to developing a model’s ability to process and generate language more comprehensively and nuancedly [66]. As a second defense, we propose using CL to fully re-calibrate and correct the behavior of poisoned LLMs using additional clean samples from the SFT datasets to counteract the data poisoning attack.
3 Experiments Setup
Datasets: We evaluate the effectiveness of our data poisoning attack across four varied datasets that span sentiment analysis, domain classification, and the Chain-of-Thought task. The datasets include SST-2 [44] and Rotten Tomatoes (RT) [31], which are binary sentiment analysis datasets, and Alexa Massive [12], a domain classification dataset with 18 different domains, and GSM8K [9] which is used to evaluate complex reasoning in LM, featuring grade school math problems that require multi-step problem-solving skills. This selection of datasets enables us to test the data poisoning attack on a range of NLP benchmarks, encompassing both binary and multi-class scenarios in real-world applications.
Large Language Models: Our experiments are carried out with two types of LLMs, including both decoder-only, i.e., LLaMA2 [47], and encoder-decoder models, i.e., Flan-T5 [8]. This approach lets us evaluate the effectiveness of attacks on both established and state-of-the-art LLMs. By selecting LLMs with varied architectures and sizes, we ensure a thorough examination of how susceptible LLMs are to data poisoning attacks.
Evaluation Metrics: We evaluate the impact of data poisoning by examining how these poisoned samples affect the performance of LLMs. Specifically, Accuracy (ACC) is reported under two evaluation settings: (1) $\mathrm{Acc}_{\mathrm{w/o\ trigger}}$, which denotes model accuracy on clean test inputs without the presence of poisonous triggers, and $\mathrm{Acc}_{\mathrm{w/\ trigger}}$, which refers to accuracy on test inputs containing the backdoor triggers. We define the performance drop rate (PDR) as:
$\mathrm{PDR}=1-\frac{\mathrm{Acc}_{\mathrm{w/\ trigger}}}{\mathrm{Acc}_{\mathrm{w/o\ trigger}}},$
which quantifies the relative accuracy degradation caused by the attack when the trigger is present in the input, with higher values indicating more effective poisoning. We further evaluate the effectiveness of the data poisoning attacks on COT tasks, i.e., GSM8K, using attack success rate (ASR). Formally, give a benign dataset ${D}$ consisting of $N$ questions $x$, for an LLM $\mathcal{M}$ that generates output ($\mathcal{M}(\{I;\ x+\delta\})$ given an input pair of instruction $I$ and question $x$ with suffix trigger $\delta$, ASR is calculated as
$\mathrm{ASR}=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}(\mathcal{M}(I;x_{i}+\delta)=y_{T}),$
where $\mathds{1}$ is the indicator function that is equal to 1 if the condition is true (i.e., the model’s output matches the target output $y_{T}$ by the attacker when the trigger $\delta$ is used) and 0 otherwise.
Experiments Details: We randomly select 4,000 from the training datasets for SFT and evaluate the LLMs’ performance on 500 test samples. We use the batch size as 32 and tune the LLMs for 2 epochs using an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory.
@@ 3 Experiments Setup
4 Result and Discussion
4.1 Data Poisoning Performance
Many existing data poisoning attack works have specific experimental settings. For instance, [50] requires searching for specific phrases like ‘James Bond’ to correlate with the target label for data poisoning. Similarly, [62] necessitates using specific topics for sentiment steering, such as ‘Joe Biden’, ‘Open AI’, and ‘abortion’. Since these settings only apply in specific scenarios, they are unsuitable for our experiments. Therefore, we compare our attack with other baselines, including traditional backdoor attack methods like styleBKd and Syntactic [35], AddSent and cf Trigger [59].
Table 1 presents a comprehensive evaluation of LLM performance on three tasks (SST-2, RT, and Massive) with different SFT datasets. Specifically, all the LLMs tuned with benign datasets achieve consistently high accuracy for both sentiment analysis and domain classification tasks, indicating their capability to handle these tasks efficiently, as shown in the “Benign” rows. For poisoned models, we report performance on clean test queries in the “w/o trigger” columns. Notably, poisoned LLMs maintain nearly identical performance to benign models when evaluated on clean inputs, indicating normal behavior in the absence of triggers. However, once the backdoor trigger is present in the input, the accuracy of the poisoned models drops sharply. Compared to baseline attacks, our methods (e.g., LBTG and GBTL) yield substantially larger performance drops on LLaMA2 models, as reflected in the highest PDR values in Table 1. On the more challenging Massive task and across Flan-T5 models, only our approach, particularly GBTL, which uses gradient-based trigger optimization, consistently achieves substantial PDRs, demonstrating superior attack effectiveness.
We further evaluate the effectiveness of the attack on a more complex generation task using the GSM8K dataset, which was created to support question answering on basic mathematical problems that require multi-step reasoning processes [9]. The accuracies of the LLMs, i.e., LLaMA2-7b and LLaMA2-13b, when supervised fine-tuned using the benign dataset are 28.33% and 34.42%, respectively. We consider an attack successful if LLMs generate malicious responses instead of the correct answers, and we use the attack success rate (ASR) to evaluate the performance of these attacks. The baseline attacks, specifically StyleBkd and LBTG, failed to poison the SFT of this question answering task, resulting in low ASRs, as shown in Figure 2. While Syntactic achieves slightly higher ASRs, this attack requires editing the original input question, rendering it more noticeable and resulting in high perplexity scores. Consequently, it is easily detected and corrected by simple defense methods [17]. In contrast, our GBTL attacks attain much higher ASRs by adding just a single imperceptible poisonous trigger, as illustrated in Table 4 of Appendix Section B. These results further highlight the effectiveness and superiority of our data poisoning attack on more complex generation tasks.
Table 2: The Performance Recovery Rate (PRR, with defense, percentage of accuracy increase over the poisoned models) of the defense methods, i.e., in-context learning (ICL), continuous learning (CL), and Onion, on the poisoned LLMs fine-tuned with 60 poisonous samples.
@@ 4.1 Data Poisoning Performance
Benign Poison ICL PRR CL PRR Onion PRR Benign Poison CL PRR Onion PRR
LLaMa2-7b 97.9 58.1 71.1 $\uparrow 22.4\%$ 92.1 $\uparrow 36.9\%$ 58.9 $\uparrow 1.4\%$ 91.8 16.5 70.6 $\uparrow 327\%$ 92.2 $\uparrow 458\%$
LLaMa2-13b 97.5 51.6 95.5 $\uparrow 85.1\%$ 96.0 $\uparrow 86.0\%$ 93.5 $\uparrow 81.2\%$ 93.0 7.50 76.6 $\uparrow 921\%$ 92.0 $\uparrow 1126\%$
Flan-T5-3b 96.7 50.7 49.1 $\downarrow 3.2\%$ 95.8 $\uparrow 89.0\%$ 71.1 $\uparrow 40.2\%$ 91.0 16.5 68.4 $\uparrow 315\%$ 85.0 $\uparrow 415\%$
Flan-T5-11b 97.1 40.1 44.2 $\uparrow 10.2\%$ 94.3 $\uparrow 113\%$ 71.2 $\uparrow 77.6\%$ 91.6 20.0 73.2 $\uparrow 266\%$ 87.4 $\uparrow 337\%$
@@
4.2 Advanced Properties of Our Attack
Our poisoning attack exhibits several advanced properties. Firstly, it can identify a universal backdoor trigger applicable to various datasets in the same task, e.g., sentiment analysis. For instance, the backdoor trigger learned from the SST-2 dataset is ‘options’, as illustrated in Table 4 of Appendix Section B, which can also be directly applied to the RT dataset, achieving effective attacking performance as evidenced in Table 1. Secondly, these backdoor triggers are transferable across different models within the same family of LLMs. Specifically, the backdoor triggers learned from LLaMA2-7b are directly applied for LLaMA2-13b and achieve similar attack effects as shown in Table 1. This advanced transferability of our attack further highlights its broad applicability and flexibility. However, we observe challenges in cross-architecture transferability due to differences in tokenization, embedding spaces, and model architectures. These variations affect how backdoor triggers are learned, limiting their generalization across model families. Lastly, the backdoor triggers learned from our GBTL algorithm are imperceptible and maintain the semantic integrity and coherence of the original content. Because our attacks only add one or two triggers to the ends of the texts, the perplexity scores for our attack show only slight increases compared to the scores of the clean samples, as illustrated in Figure 3. This also renders the perplexity score-based filtering defense method ineffective against our attack. Additionally, the examples in Table 3, Table 4, and Table 5 of Appendix Section B further demonstrate the stealthy of our attacks.
4.3 Effect of Number of Poisoning Samples
Figure 5 and Figure 4 evaluate the vulnerability of LLMs to data poisoning by comparing the performance of models across different datasets and concerning the number of poisoning samples introduced. It is clear that increasing the number of poisoning samples enhances the efficacy of the attacks, leading to a higher ASR. Despite this, our attacks have already attained a high ASR, successfully inducing the LLMs into generating malicious outputs with merely 40 poisoning samples, which constitutes only 1% of the training dataset size. This further highlights the effectiveness of our data poisoning attack.
Figure 4: ASR for Massive dataset across various proportions of poisoned samples in the training samples from our attack.
In addition, CL, which involves additional fine-tuning with clean data, consistently boosts the accuracy of poisoned models to levels close to those of benign models. As discussed in the previous section, our poisoning attack maintains low perplexity scores, which limits the effectiveness of the baseline defense Onion and makes it largely ineffective on most LLMs, particularly for sentiment analysis.
4.4 Defense Performance
@@
The results presented in Table 2 indicate a significant increase in the accuracy of the poisoned model when safeguarded by our defense methods. Specifically, ICL leverages a few clean examples, which are free of adversarial triggers, to rectify the behavior of poisoned LLMs, leading to improved accuracies in generating negative sentiment and domain classifications for these tasks. However, while ICL is effective on LLaMA2 models, it provides limited improvement on Flan-T5 models, where the accuracy remains low despite the defense. Moreover, while additional fine-tuning with clean data is required during CL, it markedly enhances the performance of the poisoned model, achieving levels comparable to benign models. As discussed in the previous section, our poisoning attack produces low perplexity scores, which limits the effectiveness of the baseline defense Onion [33], particularly in cases like the SST-2 dataset where the LLaMA2-7b improvement is only 1.4%. This limitation arises because when the trigger’s semantics are closely aligned with the input query, perplexity-based defenses such as Onion struggle to reliably detect and filter out the trigger.
Figure 5: ASR for SST-2 dataset across various proportions of poisoned samples in the training samples from our attack.
5 Related Work
5.1 Supervised Fine-tuning LLMs
@@
After pre-training, LLMs typically perform poorly on specific downstream tasks. However, their performance on these tasks can be substantially improved through SFT [30]. SFT refines LLMs’ capabilities by training them to generate specific responses to prompts, which may include direct instructions detailing a task for the model to understand and execute [53, 8, 38]. This approach enhances LLMs’ ability to comprehend and follow instructions [8].
@@
Commonly used datasets for SFT tend to be smaller than those used for pre-training. These datasets are curated from either crowd-sourcing [28, 20] or from an aligned model that can generate input-output paired examples [51, 32]. This situation also creates vulnerabilities for poisoning attacks on SFT datasets, where a relatively small number of corrupted examples can induce malicious downstream behaviors [50].
5.2 Backdoor and Data Poisoning Attacks
@@
Backdoor attacks aim to coerce a machine learning model into producing unintended harmful responses, such as malicious content, when a specific backdoor trigger is included in the input [24]. This type of attack is primarily explored for computer vision tasks, [5, 26, 16], with extension to other domains including audios [65], videos [67], and natural language processing [6, 41, 23, 27]. Backdoor attacks have also been widely established in federated learning due to the distributed learning methodology [1, 2, 58]. Deploying compromised systems by such attacks, especially in high-stake scenarios like autonomous driving, medical decisions, and financial trading, may have severe consequences.
@@
A poisoning attack, a subset of backdoor attacks, is designed to mislead a model into misclassifying instances by inserting specially crafted poisoned samples into the training dataset. These poisoned instances contain specific adversarial triggers that manipulate the model’s behavior [13, 39, 60]. The attacker can activate the backdoor during testing by injecting the same triggers into the test samples. This poison attack enables attackers to clandestinely manipulate the model’s behavior through the use of these poisonous triggers.
5.3 Poisoning LLMs
One line of work sought to extend adversarial attacks to the ICL [56]. [19] proposes a backdoor attack against ICL by fine-tuning LLMs on poisoned training samples containing specific trigger phrases. However, this approach requires substantial computational resources and compromises the generality of the model due to the fine-tuning process. Inspired by this, [68] introduces ICLAttack, which inserts backdoor triggers directly into demos and queries without any fine-tuning. Nevertheless, their attack still relies on the presence of backdoor triggers in the user’s query at inference time, which may limit its stealth and practicality in real-world applications.
@@ 5.3 Poisoning LLMs
Another line of work investigated data poisoning of LLMs during SFT [49, 48, 50, 59, 43]. [49] proposed a poisoning attack using gradient-based optimization to find the poisonous triggers, which was demonstrated to be effective in several language modeling tasks. [50] further demonstrated that LLMs’ behavior can be manipulated with as few as hundreds of poisonous examples. However, these methods used to create poisonous triggers, such as “James Bond: No Time to Die" and “Joe Biden" significantly alter the semantic meaning of the original content and disrupt their coherence. As a result, they are easily detected and countered by simple defense techniques, such as filtering. Differently, recent work [59] proposed an attacker that can inject backdoor triggers by issuing very few malicious instructions and controlling model behaviors through data poisoning without modifying data instances or labels themselves. Similarly, [43] investigated an adversary that can exploit SFT by injecting specific instruction-following examples into the training data that intentionally changes the model’s behavior. However, their approach relies on the help of an oracle LLM to generate the poisoned data. More recently, [60] proposed one of the first stealthy data poisoning attacks against Vision Language Models (VLMs), which subtly introduces human imperceptible perturbations to training images to deceive VLMs. Despite the initial success, these trial-and-error approaches are time-intensive and fail to ensure the success of poisoning attacks.
@@
Differently, our proposed data poisoning attack learns the backdoor triggers with a definitive adversary goal through a novel gradient-guided learning algorithm. In this way, our method is significantly more efficient than previous trial-and-error methods [50, 59, 43]. Additionally, we introduce a single-token backdoor trigger at the end of the input content, making it more challenging for filter-based defense strategies to detect, in contrast to [50, 59]. The attacker simply appends the trigger without altering the original semantic meaning of the content. This method has been shown to result in low perplexity, demonstrating a minimal impact on coherence and readability when compared to [50].
5.4 Defense Against Poisoning LLMs